# On the Identification of Temporally Causal Representation with Instantaneous Dependence

### v2

@@ 6 Conclusion
(Component-wise Identification of Latent Variables with instantaneous dependencies.) Suppose that the observations are generated by Equation 1 and 2, and $\mathcal{M}_{{\mathbf{c}}_{t}}$ is the Markov network over ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-1},{\mathbf{z}}_{t},{\mathbf{z}}_{t+1}\}$. Except for the assumptions A1 and A2 from Theorem 1, we further make the following assumption:
  • A3 (Sparse Latent Process): For any $z_{t,i}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{t,i}$ is an empty set.
When the observational equivalence is achieved with the minimal number of edges of estimated Markov network of $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$, then we have the following two statements:
@@
First, we demonstrate that there always exists a row permutation for each invertible matrix such that the permuted diagonal entries are non-zero [36]. By contradiction, if the product of the diagonal entry of an invertible matrix $A$ is zero for every row permutation, then we have Equation
$\text{det}(A)=\sum_{\sigma\in\mathcal{S}_{n}}\Big{(}\text{sgn}(\sigma)\prod_{i=1}^{n}a_{\sigma(i),i}\Big{)},$ (29)
@@
$\prod_{i=1}^{n}a_{\sigma(i),i}=0,\quad\forall\sigma\in\mathcal{S}_{n},$ (30)
@@
which indicates that $det(A)=0$ and $A$ is non-invertible. It contradicts the assumption that $A$ is invertible, and a row permutation where the permuted diagonal entries are non-zero must exist. Since $h_{z}$ is invertible, there exists a permuted version of the estimated latent variables, denoted as $\hat{{\mathbf{z}}}_{t}^{\pi}$, such that
$\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi(i)}}\neq 0,\quad i=1,\cdots,n.$ (31)
Since ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-1},{\mathbf{z}}_{t},{\mathbf{z}}_{t+1}\}$, by applying $\pi$ on each timestamp, we have $\pi^{\prime}$ such that
$\frac{\partial c_{t,i}}{\partial\hat{c}_{t,\pi^{\prime}(i)}}\neq 0,\quad i=1,\cdots,3n.$ (32)
@@
Second, we demonstrate that $\mathcal{M}_{{\mathbf{c}}_{t}}$ is identical to $\mathcal{M}_{\hat{\mathbf{c}}_{t}^{\pi^{\prime}}}$, where $\mathcal{M}_{\hat{\mathbf{c}}_{t}^{\pi^{\prime}}}$ denotes the Markov network of the permuted version of ${\mathbf{c}}_{t}^{\pi^{\prime}}$.
@@
On the one hand, for any pair of $(i,j)$ such that $c_{t,i},c_{t,j}$ are adjacent in $\mathcal{M}_{{\mathbf{c}}_{t}}$ while $c_{t,\pi^{\prime}(i)},c_{t,\pi^{\prime}(j)}$ are not adjacent in $\mathcal{M}_{\hat{\mathbf{c}}_{t}^{\pi^{\prime}}}$, according to Equation 26, we have $\frac{\partial c_{t,i}}{\partial c_{t,\pi^{\prime}(i)}}\cdot\frac{\partial c_{t,j}}{\partial c_{t,\pi^{\prime}(j)}}=0$, which is a contradiction with how $\pi^{\prime}$ is constructed. Thus, any edge presents in $\mathcal{M}_{{\mathbf{c}}_{t}}$ must exist in $\mathcal{M}_{\hat{\mathbf{c}}_{t}^{\pi^{\prime}}}$. On the other hand, since observational equivalence can be achieved by the true latent process $(g,f,p_{{\mathbf{c}}_{t}})$, the true latent process is clearly the solution with minimal edges.
@@
Under the sparsity constraint on the edges of $\mathcal{M}_{\hat{\mathbf{c}}_{t}^{\pi^{\prime}}}$, the permuted estimated Markov network $\mathcal{M}_{\hat{\mathbf{c}}_{t}^{\pi^{\prime}}}$ must be identical to the true Markov network $\mathcal{M}_{{\mathbf{c}}_{t}}$. Thus, we claim that
Sequentially, we further give the proof that under the same permutation ${\mathbf{z}}_{t}^{\pi}$ (which corresponds to ${\mathbf{c}}_{t}^{\pi^{\prime}}$), $z_{t,i}$ is only the function of $z_{t,\pi(i)}$. Since the permutation happens on each timestamp respectively, the cross-timestamp disentanglement is prevented clearly.
@@
Now let us focus on instantaneous disentanglement. Suppose there exists a pair of indices $i,j\in[1,\cdots,n]$. According to Equation 31, we have $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi(i)}}=0$ and $\frac{\partial z_{t,j}}{\partial\hat{z}_{t,\pi(j)}}=0$. Let us discuss it case by case.
@@
  • If $z_{t,i}$ is not adjacent to $z_{t,j}$, we have $\hat{z}_{t,\pi(i)}$ is not adjacent to $\hat{z}_{t,\pi(j)}$ as well according to the conclusion of identical Markov network. Using Equation 25, we have $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi(i)}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi(j)}}=0$, which leads to $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi(j)}}=0$.
@@
  • If $z_{t,i}$ is adjacent to $z_{t,j}$, we have $\hat{z}_{t,\pi(i)}$ is adjacent to $\hat{z}_{t,\pi(j)}$. When the Latent Process Sparsity assumption is assured, i.e., the intimate neighbor set of $z_{t,i}$ is empty, there exists at least one index $k$ such that $z_{t,k}$ is adjacent to $z_{t,i}$ but not adjacent to $z_{t,j}$. Similarly, we have the same structure on the estimated Markov network, which means that $\hat{z}_{t,\pi(k)}$ is adjacent to $\hat{z}_{t,\pi(i)}$ but not adjacent to $\hat{z}_{t,\pi(j)}$. Using Equation 26 we have $\frac{\partial z_{t,k}}{\partial\hat{z}_{t,\pi(k)}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi(j)}}=0$, which leads to $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi(j)}}=0$.
In conclusion, we always have $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi(j)}}=0$. Thus, we have reached the conclusion that
A2.4 General Case for Component-wise Identifications
@@
(General Case for Component-wise Identification.) Suppose that the observations are generated by Equation 1 and 2, and there exists ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-a},\cdots,{\mathbf{z}}_{t},\cdots,{\mathbf{z}}_{t+b}\}$ with the corresponding Markov network $\mathcal{M}_{{\mathbf{c}}_{t}}$. Suppose assumptions A1 and A2 hold true, and for any $z_{t,i}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{t,i}$ is an empty set. When the observational equivalence is achieved with the minimal number of edges of estimated Markov network of $\mathcal{M}_{\hat{{\mathbf{c}}}}$, there must exist a permutation $\pi$ of the estimated latent variables, such that $z_{t,i}$ and $\hat{z}_{t,\pi(i)}$ is one-to-one corresponding, i.e., $z_{t,i}$ is component-wise identifiable.
@@
Appendix A3 Implementation Details
A3.1 Prior Likelihood Derivation

### v3

@@ B.3 Identifiability of Latent Variables
(Component-wise Identification of Latent Variables with instantaneous dependencies.) Suppose that the observations are generated by Equation (1)-(2), $\mathcal{M}_{{\mathbf{c}}_{t}}$ and $\mathcal{M}_{{\mathbf{z}}_{t}}$ is the Markov network over ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-1},{\mathbf{z}}_{t}\}\in\mathbb{R}^{2n}$ and ${\mathbf{z}}_{t}$, respectively. Except for the assumptions A1 and A2 from Theorem 1, we further make the following assumption:
  • A3 (Sparse Latent Process): 1) For any $z_{it}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{it}$ is an empty set. 2) For any $z_{it}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{it}$ is not an empty set, but there exists a direct child $z_{jt}$, whose intimate neighbor set is an empty set.
@@
Step1: Identifiability of Markov networks
First, we demonstrate that there always exists a row permutation for each invertible matrix such that the permuted diagonal entries are non-zero (zhang2024causal). By contradiction, if the product of the diagonal entry of an invertible matrix $A$ is zero for every row permutation, then we have Equation
$\text{det}(A)=\sum_{\sigma\in\mathcal{S}_{n}}\Big(\text{sgn}(\sigma)\prod_{i=1}^{n}a_{\sigma(i),i}\Big),$ (A19)
@@
$\prod_{i=1}^{n}a_{\sigma(i),i}=0,\quad\forall\sigma\in\mathcal{S}_{n},$ (A20)
@@
which indicates that $det(A)=0$ and $A$ is non-invertible. It contradicts the assumption that $A$ is invertible, and a row permutation where the permuted diagonal entries are non-zero must exist. Since $h_{z}$ is invertible, for ${\mathbf{z}}_{t}$ at time step $t$, there exists a permuted version of the estimated latent variables, such that
$\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi_{t}(i)}}\neq 0,\quad i=1,\cdots,n,$ (A21)
where $\pi_{t}$ is the corresponding permutation at time step $t$.
@@
Second, we demonstrate that the instantaneous Markov network $\mathcal{M}_{{\mathbf{z}}_{t}}$ over ${\mathbf{z}}_{t}$ is identical to $\mathcal{M}_{\hat{\mathbf{z}}_{t}^{\pi^{\prime}}}$, where $\mathcal{M}_{\hat{\mathbf{z}}_{t}^{\pi^{\prime}}}$ denotes the Markov network of the permuted version of $\pi^{\prime}(\hat{{\mathbf{z}}}_{t})$.
@@
On the one hand, for any pair of $(i,j)$ such that $z_{t,i},z_{t,j}$ are adjacent in $\mathcal{M}_{{\mathbf{z}}_{t}}$ while $\hat{z}_{t,\pi^{\prime}(i)},\hat{z}_{t,\pi^{\prime}(j)}$ are not adjacent in $\mathcal{M}_{\hat{\mathbf{z}}_{t}^{\pi^{\prime}}}$, according to Equation (A17), we have $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi^{\prime}(i)}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t,\pi^{\prime}(j)}}=0$, which is a contradiction with how $\pi^{\prime}$ is constructed. Thus, any edge presents in $\mathcal{M}_{{\mathbf{z}}_{t}}$ must exist in $\mathcal{M}_{\hat{\mathbf{z}}_{t}^{\pi^{\prime}}}$. On the other hand, since observational equivalence can be achieved by the true latent process $(g,f,p_{{\mathbf{z}}_{t}})$, the true latent process is clearly the solution with minimal edges.
@@
Therefore, under the sparsity constraint on the edges of $\mathcal{M}_{\hat{\mathbf{z}}_{t}^{\pi^{\prime}}}$, the permuted estimated Markov network $\mathcal{M}_{\hat{\mathbf{z}}_{t}^{\pi^{\prime}}}$ must be identical to the true Markov network $\mathcal{M}_{{\mathbf{z}}_{t}}$.
Sequentially, we further provide that the Markov network between ${\mathbf{z}}_{t-1}$ and ${\mathbf{z}}_{t}$ is identifiable. We suppose that $z_{t-1,i}$ and $z_{t,j}$ are adjacent in the ground-truth Markov network while $\hat{z}_{t-1,i}$ and $\hat{z}_{t,j}$ are not adjacent in the estimated Markov network. Since $\hat{z}_{t-1,i}$ and $\hat{z}_{t,i}$ are adjacent, which implies that $\hat{z}_{t,i}$ and $\hat{z}_{t,j}$ are not adjacent in the estimated Markov network, it leads to a contradiction since $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi^{\prime}(i)}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t,\pi^{\prime}(j)}}=0$. Therefore, any edge present in the ground-truth Markov network over time-delayed latent variables must exist in the estimated one. Similarly, on the other hand, since observational equivalence can be achieved by the true latent process $(g,f,p_{{\mathbf{z}}_{t}})$, the true latent process is clearly the solution with minimal edges. Therefore, the Markov networks over the time-delayed latent variables are identifiable.
Finally, we claim that (i) the estimated Markov network $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$ is isomorphic to the ground-truth Markov network $\mathcal{M}_{{\mathbf{c}}_{t}}$.
Step2: Component-wise Identifiability of Latent Variables
Sequentially, under the same permutation $\pi_{t}$, we further give the proof that $z_{t,i}$ is only the function of $\hat{z}_{t,\pi_{t}(i)}$. According to the A3, we consider two cases: 1) For any $z_{it}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{t,i}$ is empty. 2) For any $z_{t,i}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{t,i}$ is not empty, but there exists a direct child $z_{zj}$, whose intimate neighbor is empty.
We first consider the first case, where the intimate neighbor set of $z_{t,i}$ is an empty set. Since the permutation happens on each time step, the cross-time disentanglement is prevented clearly.
@@
Now, let us focus on instantaneous disentanglement. Suppose there exists a pair of indices $i,j\in\{1,\cdots,n\}$. According to Equation (A21), we have $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi_{t}(i)}}=0$ and $\frac{\partial z_{t,j}}{\partial\hat{z}_{t,\pi_{t}(j)}}=0$. Let us discuss it case by case.
@@
  • If $z_{t,i}$ is not adjacent to $z_{t,j}$, we have $\hat{z}_{t,\pi_{t}(i)}$ is not adjacent to $\hat{z}_{t,\pi_{t}(j)}$ as well according to the conclusion of identical Markov network. Using Equation (A14), we have $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi_{t}(i)}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi_{t}(j)}}=0$, which leads to $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi_{t}(j)}}=0$.
@@
  • If $z_{t,i}$ is adjacent to $z_{t,j}$, we have $\hat{z}_{t,\pi_{t}(i)}$ is adjacent to $\hat{z}_{t,\pi_{t}(j)}$. When the Assumption A3 (Sparse Latent Process) is assured, i.e., the intimate neighbor set of $z_{t,i}$ is empty, there exists at least one pair of $(t^{\prime},k)$ such that $z_{t^{\prime},k}$ is adjacent to $z_{t,i}$ but not adjacent to $z_{t,j}$. Similarly, we have the same structure on the estimated Markov network, which means that $\hat{z}_{t^{\prime},\pi_{t^{\prime}}(k)}$ is adjacent to $\hat{z}_{t,\pi_{t}(i)}$ but not adjacent to $\hat{z}_{t,\pi_{t}(j)}$. Using Equation (A17) we have $\frac{\partial z_{t,k}}{\partial\hat{z}_{t^{\prime},\pi_{t^{\prime}}(k)}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi_{t}(j)}}=0$, which leads to $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi_{t}(j)}}=0$.
Sequentially, we further consider the second case, where the intimate neighbor set of $z_{t,i}\in{\mathbf{z}}_{t}$ is not empty, but there exists a direct child $z_{t,j}$, whose intimate neighbor set is an empty set. According to Theorem A1, the second case means that there exists an estimated latent variable $\hat{z}_{t,k}$ and a function $h_{k}$, such that $\hat{z}_{t,k}=h_{k}(z_{t,i},z_{t,j})$.
The second case denotes that $z_{t,i}$ is independent of the remaining latent variables given ${\mathbf{z}}_{t-1}$ and $z_{t,j}$, which can be formalized as:
$\log p(z_{t,i},z_{t,j},{\mathbf{z}}_{t,\\ /i,j}|{\mathbf{z}}_{t-1})=\log p(z_{t,i}|{\mathbf{z}}_{t-1})+\log p(z_{t,j}|{\mathbf{z}}_{t-1},z_{t,i})+\log p({\mathbf{z}}_{t,\\ /i,j}|{\mathbf{z}}_{t-1}).$ (A22)
Since the latent structures of ${\mathbf{z}}_{t-1}$ and ${\mathbf{z}}_{t}$ are the same, the aforementioned equation can be further rewritten as:
$\log p(z_{t,i},z_{t,j},{\mathbf{z}}_{t,\\ /i,j}|{\mathbf{z}}_{t-1})=\log p({\mathbf{z}}_{t,i}|{\mathbf{z}}_{t-1,i})+\log p({\mathbf{z}}_{t,j}|{\mathbf{z}}_{t-1,j},{\mathbf{z}}_{t,i})+\log p({\mathbf{z}}_{t,\\ /i,j}|{\mathbf{z}}_{t-1,\\ /i,j}).$ (A23)
According to Equation (A6), we can further have:
[…]
$=$ $\log p(z_{t,i}|z_{t-1,i})+\log p(z_{t,j}|z_{t-1,j},z_{t,i})+\log p({\mathbf{z}}_{t,\backslash i,j}|{\mathbf{z}}_{t-1,\backslash i,j})+\log|\mathbf{J}_{h_{z},t}|.$
Sequentially, we further conduct the first-order derivative w.r.t $\hat{z}_{t,\pi(j)}$, so the aforementioned equation can be rewritten as follows:
[…]
Since the intimate neighbor set of $z_{t,j}$ is empty, implying that $z_{t,j}$ is component-wise identifiable according to case (1), so $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,\pi(j)}}=0$ and $\frac{\partial z_{t,l}}{\partial\hat{z}_{t,\pi(j)}}=0$. As a result, Equation (A25) can be written as:
$\frac{\partial\log p(\hat{z}_{t,\pi(j)}|\hat{z}_{t-1,\pi(j)},\hat{z}_{t,k})}{\partial\hat{z}_{t,\pi(j)}}=\frac{\partial\log p(z_{t,j}|z_{t-1,j},z_{t,i})}{\partial z_{t,j}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t,\pi(j)}}+\frac{\partial|\mathbf{J}_{h_{z},t}|}{\partial\hat{z}_{t,\pi(j)}}.$ (A26)
Moreover, we further conduct the second-order derivative w.r.t $\hat{z}_{t-1,k}$, then we have:
$0=\frac{\partial^{2}\log p(z_{t,j}|z_{t-1,j},z_{t,i})}{(\partial z_{t,j})^{2}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t,\pi(j)}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t-1,k}},$ (A27)
which means that $\frac{\partial z_{t,j}}{\partial\hat{z}_{t,\pi(j)}}\cdot\frac{\partial z_{t,j}}{\hat{z}_{t-1,k}}=0$ with the sufficient variability of $\frac{\partial^{2}\log p(z_{t,j}|z_{t-1,j},z_{t,i})}{(\partial z_{t,j})^{2}}$. Since $\frac{\partial z_{t,j}}{\partial\hat{z}_{t,\pi(j)}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t-1,k}}=\frac{\partial z_{t,j}}{\partial\hat{z}_{t,\pi(j)}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial\hat{z}_{t,k}}{{\partial\hat{z}_{t-1,k}}}$. Moreover, $\frac{\partial z_{t,j}}{\partial\hat{z}_{t,\pi(j)}}\neq 0$ and $\frac{\partial\hat{z}_{t,k}}{\partial\hat{z}_{t-1,k}}\neq 0$, then we have $\frac{\partial z_{t,j}}{\partial\hat{z}_{t,k}}=0$, implying that $z_{t,j}$ is not the function of $\hat{z}_{t,k}$. Since we have $\hat{z}_{t,k}=h_{k}(z_{t,i},z_{t,j})$, then we have $\hat{z}_{t,k}=h_{k}(z_{t,i})$, implying that $z_{t,i}$ is component-wise identifiable.
Thus, we have reached the conclusion that (ii) there exists a permutation $\pi$ of the estimated latent variables, such that $z_{t,i}$ and $\hat{z}_{t,\pi(i)}$ is one-to-one corresponding, i.e., $z_{t,i}$ is component-wise identifiable.
B.4 General Case for Component-wise Identifications
@@
(General Case for Component-wise Identification.) Suppose that the observations are generated by Equation (1)-(2), and there exists ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-a},\cdots,{\mathbf{z}}_{t},\cdots,{\mathbf{z}}_{t+b}\}\in\mathbb{R}^{(a+b+1)\times n}$ with the corresponding Markov network $\mathcal{M}_{{\mathbf{c}}_{t}}$. Suppose assumptions A1 and A2 hold true, and for any $z_{t,i}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{t,i}$ is an empty set. When the observational equivalence is achieved with the minimal number of edges of estimated Markov network of $\mathcal{M}_{\hat{{\mathbf{c}}}}$, there must exist a permutation $\pi$ of the estimated latent variables, such that $z_{t,i}$ and $\hat{z}_{t,\pi(i)}$ is one-to-one corresponding, i.e., $z_{t,i}$ is component-wise identifiable.
@@
Figure A5: An example of DAG (a) and Markov Network (b).
B.5 Identifiability of Latent Causal Process
Building on the results of Theorem 2, the latent variables are component-wise identifiable, which directly implies the identifiability of the latent causal process up to the Markov equivalence class. Leveraging temporal information allows one to further refine this identification beyond the equivalence class. If each latent variable $z_{t}$ has at least one temporal ancestor, one can establish full identifiability of the graph over the latent processes.
Theorem A3.
(Identification of Latent Causal Process.) Suppose that the observations are generated by Equation (1)-(2), and that $\mathcal{M}_{{\mathbf{c}}_{t}}$ is the Markov network over ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-1},{\mathbf{z}}_{t}\}\in\mathbb{R}^{2n}$. Suppose that all assumptions for Theorem 2 hold. We further make the following assumption: for any pair of adjacent latent variables $z_{t,i},z_{t,j}$ at time step $t$, their time-delayed parents are not identical, i.e., $\text{Pa}_{d}(z_{t,i})\not=\text{Pa}_{d}(z_{t,j})$. Then the causal graph of the latent causal process is identifiable.
Proof.
Since all assumptions for Theorem 2 hold, latent variable ${\mathbf{z}}_{t}$ is component-wise identifiable, i.e., there exists a permutation $\pi$ and invertible functions $h_{i}$ such that $z_{t,i}=h_{i}(\hat{z}_{t,\pi(i)})$. As ${\mathbf{c}}_{t}$ is nothing but a concatenation of ${\mathbf{z}}_{t}$ from different time steps, ${\mathbf{c}}_{t}$ is also component-wise identifiable with corresponding $\pi^{\prime}$ and $h_{i}^{\prime}$. Therefore, their conditional independence relationships are consistent:
$c_{t,i}\perp c_{t,j}\mid\{c_{t,k}\mid k\in S\}\Rightarrow\hat{c}_{t,\pi^{\prime}(i)}\perp\hat{c}_{t,\pi^{\prime}(j)}\mid\{\hat{c}_{t,\pi^{\prime}(k)}\mid k\in S\}$ (A28)
[…]
We first prove that the skeleton of the estimated causal graph is identical to the true skeleton.
  • On the one hand, if $c_{t,i}$ and $c_{t,j}$ are not adjacent, there exists a d-separation set $S_{d}$ such that $c_{t,i}\perp c_{t,j}\mid\{c_{t,k}\mid k\in S_{d}\}$. Meanwhile, we have $\hat{c}_{t,\pi^{\prime}(i)}\perp\hat{c}_{t,\pi^{\prime}(j)}\mid\{\hat{c}_{t,\pi^{\prime}(k)}\mid k\in S_{d}\}$. Thus, $\hat{c}_{t,\pi^{\prime}(i)}$ and $\hat{c}_{t,\pi^{\prime}(j)}$ are not adjacent as well.
  • On the other hand, if $c_{t,i}$ and $c_{t,j}$ are adjacent, there will be no d-separation set for $\hat{c}_{t,\pi^{\prime}(i)},\hat{c}_{t,\pi^{\prime}(j)}$, meaning that $\hat{c}_{t,\pi^{\prime}(i)}$ and $\hat{c}_{t,\pi^{\prime}(j)}$ are adjacent in this case.
In conclusion, the skeleton is identifiable.
Then we prove that the time-delayed edges can be identified. Consider any $c_{t,i}$ from ${\mathbf{z}}_{t-1}$ and $c_{t,j}$ from ${\mathbf{z}}_{t}$. Similarly, we have $\hat{c}_{t,\pi^{\prime}(i)}$ from $\hat{{\mathbf{z}}}_{t-1}$ and $\hat{c}_{t,\pi^{\prime}(j)}$ from $\hat{{\mathbf{z}}}_{t}$. Since the skeleton is identified and the direction is implied by temporal information, specifically going from time step t1t-1 to $t$, there exists an edge from $\hat{c}_{t,\pi^{\prime}(i)}$ to $\hat{c}_{t,\pi^{\prime}(j)}$ if and only if $c_{t,i}$ points to $c_{t,j}$. Thus the temporally latent causal process can be identified.
Finally, we determine the direction of instantaneous edges. Since the skeleton is already identifiable, we only need to determine the direction. Consider any pair of $c_{t,i}$ and $c_{t,j}$ from ${\mathbf{z}}_{t}$, where $c_{t,i}\rightarrow c_{t,j}$. Since $\text{Pa}_{d}(c_{t,i})\not=\text{Pa}_{d}(c_{t,j})$, there exists at least one $c_{t,k}$ from ${\mathbf{z}}_{t-1}$ such that $c_{t,k}$ points to exactly one of $c_{t,i}$ or $c_{t,j}$. We proceed with a case-by-case analysis.
  • In case $1$, we have $c_{t,k}\rightarrow c_{t,i}\rightarrow c_{t,j}$ where $c_{t,k},c_{t,j}$ are not adjacent, and $c_{t,i}$ is in the d-seperation of $c_{t,k}$ and $c_{t,j}$. Meanwhile, in the estimated skeleton, $\hat{c}_{t,\pi^{\prime}(i)}$ is also in the d-seperation of $\hat{c}_{t,\pi^{\prime}(k)}$ and $\hat{c}_{t,\pi^{\prime}(j)}$. Since the direction of time-delayed edge $\hat{c}_{t,\pi^{\prime}(k)}\rightarrow\hat{c}_{t,\pi^{\prime}(i)}$ is the known in the estimated skeleton $\hat{c}_{t,\pi^{\prime}(k)}-\hat{c}_{t,\pi^{\prime}(i)}-\hat{c}_{t,\pi^{\prime}(j)}$, we deduce $\hat{c}_{t,\pi^{\prime}(i)}\rightarrow\hat{c}_{t,\pi^{\prime}(j)}$ based on the d-seperation.
  • In case $2$, we have $c_{t,i}\rightarrow c_{t,j}\leftarrow c_{t,k}$ where $c_{t,k},c_{t,i}$ are not adjacent. Similarly, we have that $\hat{c}_{t,\pi^{\prime}(j)}$ is not in the d-seperation of $\hat{c}_{t,\pi^{\prime}(k)}$ and $\hat{c}_{t,\pi^{\prime}(i)}$, and it can be shown that $\hat{c}_{t,\pi^{\prime}(i)}\rightarrow\hat{c}_{t,\pi^{\prime}(j)}$.
In conclusion, the identifiability of the instantaneous latent causal graph is established. Consequently, the entire latent causal process, encompassing both the time-delayed and instantaneous components, is identifiable up to the causal graph.
Discussion: To further illustrate this Corollary, we use dataset A as an example, whose causal graph is shown in Figure A6. Since the skeleton and directions of time-delayed edges are straightforward to determine, we primarily focus on analyzing the directions of instantaneous edges within ${\mathbf{z}}_{t}$. Since $z_{t-1,3}\rightarrow z_{t,3}\leftarrow z_{t,2}$ is a v-structure, $z_{t,3}\leftarrow z_{t,2}$ can be determined. Since $z_{t-1,1}\rightarrow z_{t,1}\rightarrow z_{t,2}$ is a chain, $z_{t-1,1}$ and $z_{t,2}$ are not adjacent, and $z_{t-1,1}\rightarrow z_{t,1}$ is known, we have $z_{t,1}\rightarrow z_{t,2}$. Thus, the causal graph is identifiable. As shown in Figure 4, the model can learn the true causal graph.
Figure A6: The example of dataset A whose causal graph is identifiable.
B.6 Markov Network
A Markov network (or Markov random field) is a graphical model that represents the joint distribution of a set of random variables using an undirected graph.
Definition 3 (Markov Network).
Markov network is an undirected graph $G=(V,E)$ with a set of random variables $X_{v\in V}$, where any two non-adjacent variables are conditionally independent given all other variables. That is,
$X_{a}\perp X_{b}|X_{V\backslash\{a,b\}},\quad\forall(a,b)\notin E.$ (A29)
Markov Networks and Directed Acyclic Graphs (DAGs) are both graphical models employed to represent joint distributions and to illustrate conditional independence properties. As shown in Figure A5, both of them are utilized to describe the latent causal process, yet they do not have to be equivalent.
B.7 Isomorphism of Markov networks
Definition 4 (Isomorphism of Markov networks).
We let the $V(\cdot)$ be the vertical set of any graphs, an isomorphism of Markov networks $M$ and $\hat{M}$ is a bijection between the vertex sets of $M$ and $\hat{M}$
$V(M^)\color[rgb]{0,0,0}f$$\rightarrow V(\hat{M})$
such that any two vertices $u$ and $v$ of $M$ are adjacent in $G$ if and only if $f(u)$ and $f(v)$ are adjacent in $\hat{M}$.
B.8 Illustration of Intimate Neighbor Set
Here, we provide an example for a better understanding of the Intimate Neighbor Set. Take Figure A5(b) as an example. If we consider only one time step, says, $\mathbf{z}_{t}$, $\{z_{t,1},z_{t,2},z_{t,3}\}$ forms a clique. Thus we have $\Psi(z_{t,1})=\{z_{t,2},z_{t,3}\},$ since $z_{t,2}$ is adjacent to $z_{t,1}$ and all other neighbours of $z_{t,1}$, i.e., $z_{t,3}$. Similarily, we have $\Psi(z_{t,1})=\{z_{t,1},z_{t,3}\},\Psi(z_{t,1})=\{z_{t,1},z_{t,2}\}$ as well. In this case, none of them is identifiable. In contrast, if we take all 3 time steps into consideration, the Intimate set of all latent varibles becomes empty. For example, $z_{t,2}$ is adjacent to $z_{t,1}$ but not adjacent to at least one neighbour of $z_{t,1}$, i.e., $z_{t-1,1},$ thus $z_{t,2}\not\in\Psi(z_{t,1}).$ Similarly, do this test for all other pairs variables, and the conclusion can be achieved.
Appendix C Discussion of Assumptions
To enhance understanding of our theoretical results, we provide some explanations of the assumptions, their connections to real-world scenarios, as well as the potential boundary of theoretical results.
First, The smooth and positive density assumption is standard in the literature on nonlinear ICA (khemakhem2020variational; yao2022temporally; kong2022partial), meaning that the latent variables ${\mathbf{z}}_{t}$ change continuously based on historical information. For instance, in a weather dataset, temperature varies smoothly over time. However, this assumption may be violated if we cannot fully capture the transition probabilities from the observations. To mitigate this, we can sample a larger dataset to better estimate the latent causal process.
Second, the sufficient variability assumption has also been commonly adopted for the existing identifiable results of temporally causal representation learning (yao2021learning; yao2022temporally; chen2024caring). This assumption describes the changeability of latent variables. Take human motion forecasting as an example, the latent variable may represent joint location at different time steps. The linear independence of the latent variables means that the changes of each joint cannot be linearly represented by others. Besides, while the sufficiency assumption is foundational to the theory of identifiability, it is not excessively restrictive. Even in cases where this assumption is not fully satisfied, it is still possible to achieve a degree of subspace identifiability (kong2022partial).
Finally, the sparse latent process is the key assumption of our method, which is common in real-world scenarios. Intuitively, in human motion forecasting, the joints of humans can be considered as latent variables. Since there are few connections among the joints of the human body, the latent process described by the skeleton motion trajectory is sparse. Even though the sparsity assumption is not completely satisfied, we can still attain a subspace level of identifiability (li2024subspace). In this case, each true variable can be a function of, at most, an estimated version of its corresponding variable and those within the intimate set. Let us provide a simple example here. In a video of a moving car, it might be hard to have individual identifiability of the separate car wheels and car body; however, they can be considered as essential parts of the macro variable ’car’. This macro representation might be sufficient for the purpose of modeling the interactions between the car and other objects.
Appendix D The importance of considering instantaneous dependence
Here, we provide an example as shown in Figure A7 to show why instantaneous dependence is important in time-series modeling. As shown in Figure A7 (a), without considering the instantaneous dependencies of a knee and an ankle, an abnormal motion might be generated, where the leg is bent at a distorted angle. Meanwhile, by taking the instantaneous dependencies into account as shown in Figure A7 (b), the predicted motion complies with human physiological structure.
Figure A7: An example of human motion forecasting, where joints can be considered as latent variables and the latent skeleton decides the motion.
Appendix E Implementation Details
E.1 Prior Likelihood Derivation